OpenStudy (anonymous):

Suppose that a constant force is applied to an object. Newton's Second Law of Motion states that the acceleration of the object varies inversely with its mass. A constant force acting upon an object with mass 8kg produces an acceleration of 7m/s (squared) . The acceleration of another object produced by the same force is 4m/s(squared) . What is the mass of this object?

OpenStudy (anonymous):

\[a=\frac{force}{mass}\]

OpenStudy (anonymous):

A constant force acting upon an object with mass 8kg produces an acceleration of 7m/s (squared) plug the info in formula above \[7 m/s^2=\frac{f}{8 kg}\] solve for f

OpenStudy (anonymous):

once you found f plug it in here to solve for Mass \[a=\frac{force}{mass}\] "The acceleration of another object produced by the same force is 4m/s(squared)" \[4 m/s^2=\frac{force}{mass}\]